In a recent survey of 1159 students, 968 of them would like to recommend a website to their friends. Construct a 95% confidence interval to estimate the proportion of all students who would recommend the website to their friends.

Give your answers to three decimals:


_____,_____


1
Expert's answer
2022-05-12T02:47:19-0400

The sample proportion is computed as follows


"\hat{p}=\dfrac{X}{N}=\dfrac{968}{1159}"

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{N}})"

"=(\dfrac{968}{1159}-1.96\sqrt{\dfrac{\dfrac{968}{1159}(1-\dfrac{968}{1159})}{1159}},"


"\dfrac{968}{1159}+1.96\sqrt{\dfrac{\dfrac{968}{1159}(1-\dfrac{968}{1159})}{1159}})"


"=(0.814,0.857)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.814 < p < 0.857," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.814, 0.857)."